A 10-year Treasury bond with par value of $1,000 has a 6% p.a. coupon rate and pays interest every six months. The bond is four years old and has just made its eighth payment. The market now requires a 7% p.a. return on the bond. What is the expected price of the bond?

a.

$965.63

b.

$981.63

c.

$809.34

d.

$951.68

e.

$1,000.00
Answer #1

Time to maturity is =10-4 =6 years or 12 periods

Price of the bond is:-

=PV(rate,nper,pmt,fv)

=PV(7%/2,12,6%/2*1000,1000)

=951.68